OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~\log _{3}2=0.631~\log _{3}5=1.465,find~the~values~of:\]\[\log _{3}0.1\]

OpenStudy (anonymous):

\[\log _{3}\frac{ 1 }{ 10 }\]\[\log _{3}\frac{ 1 }{ 2\times5 }\]\[\log _{3}1-(\log _{3}2+\log _{3}5)\]\[0-0.631-1.465=-2.096\]

OpenStudy (anonymous):

can we do like this? \[=\log 0.1~\div~\log3\]\[\approx-2.096\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

OpenStudy (kainui):

Well what you've done seems correct up to this point: \[=\log_3 0.1~\div~\log_33\]\[\approx-2.096\] This doesn't really say anything since \[\Large \log_33=1\] So since this is already true : \[\Large \log_3 0.1 \approx -2.096\] All you're doing is dividing by 1 on both sides, which doesn't really do anything.